OpenStudy (anonymous):

Let f(x) = 4x2 and g(x) = 7x. Determine the domain of the composite function and then find the composite function. (f g)(x)

OpenStudy (anonymous):

Domain all real numbers. put 7x into f(x) every where there is an x f(x) = 4(7x)^2 f(x) = 4*49x^2 f(x) = 196x^2

OpenStudy (anonymous):

bah all of those problems today arnt rite. i worked and worked on em but there not rite thanx tho